geerky42 (geerky42):

y = x² + 3, find dy/dt, given that dx/dt = 2 when x = 1

OpenStudy (calculusfunctions):

Here we wish to differentiate with respect to t. Do you know how?

OpenStudy (anonymous):

You can take the derivative implicitly...

OpenStudy (anonymous):

Just differentiate it with respect to t...

geerky42 (geerky42):

\[\frac{dy}{dt} = 2x\frac{dx}{dt}\] Now what?

OpenStudy (calculusfunctions):

Perfect!!

OpenStudy (calculusfunctions):

Now what?

OpenStudy (anonymous):

Now substitute the values you are given with...

OpenStudy (calculusfunctions):

Please let her answer @waterineyes ?

OpenStudy (anonymous):

Carry on @calculusfunctions

geerky42 (geerky42):

substitute? how? we only know that dx/dt = 2 WHEN x = 1

geerky42 (geerky42):

beside i dont find how i can use this information

OpenStudy (calculusfunctions):

Correct so in your last step, substitute these values and find dy/dt.

geerky42 (geerky42):

wait so question are asking me to find dy/dt when x = 1?

OpenStudy (calculusfunctions):

Yes!!

geerky42 (geerky42):

oh well it makes sense. thanks.

OpenStudy (calculusfunctions):

So then dy/dt = ? what is your answer?

geerky42 (geerky42):

\[\Huge \frac{dy}{dt} = 4\]

OpenStudy (calculusfunctions):

Beautiful!!

geerky42 (geerky42):

lol thanks